OpenStudy (anonymous):

What is the value of n in the equation 12n + 8 = 11? Please submit your answer as a reduced fraction.

OpenStudy (anonymous):

12n + 8 = 11 12n = 11 - 8 12n = 3 n= 3/12 n= 1/4

OpenStudy (anonymous):

thanks